Center of Gravity Method: East Texas Seasonings is preparing to build one processing center to serve its four sources of seasonings. The four source locations are at coordinates shown below. Also, the volume from each source is provided. What is the center of gravity? 

  X-coordinate Y-coordinate Volume
Athens, Texas 30 30 150
Beaumont, Texas 20 10 350
Carthage, Texas 10 70 100
Denton, Texas 50 50 200
Answer #1

Solution:

The center of gravity coordinates can be calculated as;

X-coordinate = Sum of [Xi x Volume] / Total Volume

Total Volume = 150 + 350 + 100 + 200 = 800

X-coordinate = [(30 x 150) + (20 x 350) + (10 x 100) + (50 x 200)] / 800

X-coordinate = 22,500 / 800

X-coordinate = 28.13

Y-coordinate = Sum of [Yi x Volume] / Total Volume

Y-coordinate = [(30 x 150) + (10 x 350) + (70 x 100) + (50 x 200)] / 800

Y-coordinate = 25,000 / 800

Y-coordinate = 31.25

Center of gravity: X = 28.13, Y = 31.25